OpenStudy (kayluh123):

The area of a rectangle is 63yd^2, and the length of the rectangle is 5yd more than twice the width. Find the dimensions of the rectangle.

OpenStudy (kayluh123):

Length and Width

OpenStudy (anonymous):

since the length is L = 2W +5, and W = W then W * (2W +5) = 63

OpenStudy (anonymous):

which simplifies to 2w^2 +5w = 63 and that to 2w^2 + 5w -63 = 0 solve for w and pick only the positive number that's your answer